HUBBLE SPACE TELESCOPE DAILY REPORT #2495 PERIOD COVERED: 0000Z (UTC) 11/05/99 - 0000Z (UTC) 11/08/99 Daily Status Report as of 312/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8252 (Extinction Curves of Distant Galaxies) The WF/PC-2 was used to make observations of the distant galaxy/quasar LBQS1009-0 in order to study its extinction. The program will more than double the number of galaxies in which extinction curves have been measured and it will provide the first such measurements far outside the Local Group. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Five Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.3 Completed Three Sets of STIS/CCD 8409 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The observations completed with no problems. 1.4 Completed Seven Sets of STIS/CCD 8549 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations as part of a non-scripted POMS Test Proposal. As described in HSTAR 7341 and the previous report, the acquisition for the first iteration of this proposal defaulted to fine lock back-up on one FGS only, possibly affecting one observation. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.5 Completed Two Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. As described in HSTAR 7341 and the previous report, the acquisition for the first iteration of this proposal defaulted to fine lock back-up on one FGS only, possibly affecting two observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.6 Completed STIS/CCD 8104 (Constraints on the Evolution of Powerful Radio Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of 3C67. Recent work has identified the GHz Peaked Spectrum and Compact Steep Spectrum radio sources as the most likely candidates for the progenitors of the large-scale powerful classical double sources. Both a wavelength calibration image and a CCD flat calibration image were also taken. The observations were completed as planned, and no anomalies were reported. 1.7 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local group galaxies offer the opportunity for a closer look at young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.8 Completed STIS/CCD/MA1/MA2 8243 (SINS: The Supernova Intensive Study - Cycle 8) The Space Telescope Imaging Spectroscope (CCD/MA1/MA2) was used to make target of opportunity observations of supernova SN1999em. Supernovae reveal the last stages of stellar evolution, create the chemical history of the Universe, energize the interstellar gas, and occupy the center of attention for the extragalactic distance scale and cosmology. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.9 Completed FOC/96 7923 (FOC F/96 Turn-on) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 8345 (Multipolar Bubbles and Jets in Very Low Excitation Planetary Nebulae - A WF/PC-2 H-Alpha Imaging Survey) The WF/PC-2 was used to make H-Alpha observations of the very low excitation planetary nebula PK068-02D1 in order to study its multipolar bubbles and jets. Comparison of the morphology of this young PNe sample with more evolved PNe populations will set new constraints on models for the production of jet-like outflows and nebular evolution. The observations were completed as planned, and no problems were reported 1.11 Completed Four Sets of WF/PC-2 8515 (WF/PC-2 Pre-SM3A UV Baseline Observations) The WF/PC-2 was used to make far-ultraviolet observations of the CVZ targets 1713+695, 1104+602, 2126-734 and 0446+789 as part of the WF/PC-2 pre-SM3A UV baseline calibration observations. The observations were completed as planned, and no problems were reported. 1.12 Completed Two Sets of STIS/CCD 8545 (POMS Test Proposal: STIS non-scripted parallel proposal continuation VI) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.13 Completed STIS/CCD 8477 (Star Formation Knots in the Distant Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to obtain long slit spectra in star-formation knots in NGC428. The observations completed with no reported problems. 1.14 Completed WF/PC-2 8403 (A Rare Case of Major LBV Eruption) The WF/PC-2 was used to follow in detail the spectrophotometric evolution of NGC2363-V1 until the end of its major eruption. The observations completed with no reported problems. 1.15 Completed Two Sets of STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ0356-5, HCLQ0250-3, and HCLQ0516-4 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed STIS/CCD/MA1/MA2 8245 (Black Hole Accretion Outbursts in Soft X- Ray Transients) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the optical and ultraviolet counterpart of a soft X-ray outburst region from a black hole accretion disk. A wavelength and a CCD flat calibration images were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed STIS/CCD 8415 (CCD Sparse-Field CTE External C8) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of M67. This program will allow the derivation of accurate coefficients that can be used to correct low count level data {both imaging and spectroscopy} for nonlinearity/CTE effects. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed STIS/CCD 8226 (Discerning the Origins of Blue Stragglers: Masses and Rotation Rates) The Space Telescope Imaging Spectrograph (CCD) was used to study the blue stragglers in globular cluster NGC-6752-BS. Conventional wisdom holds that the masses of Blue Straggler Stars in globular clusters are significantly larger than the turnoff mass. A wavelength calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed WF/PC-2 8398 (Saturn's Rings) The WF/PC-2 was used to make observations of the rings around the planet Saturn. Saturn's rings are a unique dynamical laboratory and continue to present challenging problems that require additional data to attack. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed WF/PC-2 7374 (Probing Low Density Cluster Environments at Moderate-to-High Redshifts) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-0723. Recent HST WF/PC-2 images of rich clusters at moderate-to-high redshifts have revealed that morphological information is a crucial ingredient in understanding the physical processes responsible for the evolution of the galaxy population in clusters. The observations were executed as scheduled, and no anomalies were noted. 1.21 Completed WF/PC-2 8175 (A Survey of Gravitational Lenses as Cosmological Tools III) The WF/PC-2 was used to make observations of the gravitational lens system B1933+507. Gravitational lenses are powerful tools for studying cosmology, galactic structure, and galaxy evolution. The observations were completed as planned, and no anomalies were reported. 1.22 Completed STIS/CCD/MA1 8112 (Probing the Galactic Halo and Beyond with Young Supernovae) The Space Telescope Imaging Spectroscope (CCD/MA1) was used for a target of opportunity program to obtain STIS echelle spectra of a bright new supernova {V<14} to characterize the ionization state, gas-phase abundances, depletion on dust grains, metallicity, and gas kinematics in the Galactic halo, the ISM and halo of the SN's host galaxy, and, if the properties of the sightline are favorable, in the intervening intergalactic medium. Approximately one hour into this six-and-one-half-hour proposal, the MCE-1 reset mentioned in 2.3 occurred. About an hour later, MAMA-1 was recovered and the proposal continued. At this time, the exact effect on the proposal is not known. Otherwise, the observations completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 27 Successful: 27 Scheduled Re-acquisitions: 19 Successful: 19 2.2 FHST Updates: Scheduled: 70 Successful: 70 2.3 Operations Notes: The SSR EDAC error counter was cleared nine times per ROP SR-1. The CCS "A" string went down due to a failed UPS unit. Operations transferred to "B" string at 309/1641Z. Then, real-time operations transferred to the "G" string at 310/0145Z in preparation for JIS #6 (Rendezvous/Capture) which will be conducted for approximately 12 hours beginning on Monday morning. The "G" string went down for unknown reasons at 311/2103Z, but was back on-line at 311/2139Z. The operations team moved to the back-up facility in Building 23 at 312/0900Z.. ROP SR-1 was utilized six times to re-dump important data from target of opportunity proposals (1.8 and 1.22). The SSA transmitter was commanded on and off twice per ROP IC-2. The NSSC-1 status buffer was dumped at 309/2323Z as directed by ROP NS-3 Using ROP SR-1, the TRI module was re-configured at 311/0414Z and at 311/1935Z. There was a STIS MCE-1 reset at 311/145928Z while the high voltage was on and while the HST was in an SAA interval. The target of opportunity proposal described in 1.22 was affected (see 1.22 for details). The STIS flight software error counter was reset at 311/1600Z. MAMA-1 was recovered at 311/1705Z, the next time that high voltage was commanded on by normal stored commanding. The SSR science playback pointer was set at 312/0119Z after a re-dump of science data. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW